Okay this one got me.
Carlo's age is 7/10 of mark's. Four years from now, carlo's age will be 8/11 of mark's age. How is carlo?
This is my equation.
X is carlo
7/10x is mark
X +4 = 8/11(7/10x + 4)
Am i on the right track?

Answer by rothauserc(4718)

let x be Carlo's age and y be Mark's age, then we have two equations
x = 7y/10
x + 4 = 8(y+4)/11
substitute for x in second equation
(7y/10) + 4 = 8(y+4)/11
denominator will be 10*11 = 110
11*(7y) + 4*110 = 80(y+4)
77y +440 = 80y +320
3y = 120
y = 40
x = 7*40 / 10 = 28
Carlo's age is 28 and Mark's age is 40